# GazeTrak: Exploring Acoustic-based Eye Tracking on a Glass Frame

### v1

@@ 1. Introduction
We conducted two rounds of user studies to evaluate the performance of GazeTrak. During the studies, each participant was asked to look at and follow the instruction points on the screen. In the first round of the study, 12 participants evaluated our first hardware prototype (shown in Fig. 3 (f)), where the microphones and speakers were glued on a glass frame. The average cross-session tracking accuracy was . It confirmed the optimal settings of the sensing system, which helped us design the final prototype. The final prototype (as shown in Fig. 3 (g)) features a more compact form factor, significantly lower signal strength, and improved environmental sustainability. To ensure consistent performance between the two prototypes, we conducted a second round of study with 10 participants, including some new participants, evaluating the final prototype. The final prototype achieved an average tracking accuracy of for cross-session scenarios and for in-session scenarios with a refreshing rate of 83.3 Hz. We made a demo video111https://youtu.be/XvNLNkfQY7Q to demonstrate the tracking performance and real-world applications of our system.
@@ 2.3. Wearable Eye Tracking Technologies
To overcome the challenges that non-wearable eye tracking technologies face as described in the last two subsections, many wearable eye tracking technologies based on cameras (Zhang et al., 2011; Ryan et al., 2008; Noris et al., 2011; Cho et al., 2012; Kassner et al., 2014; Lanata et al., 2015; Mayberry et al., 2014, 2015), optical sensors (Borsato and Morimoto, 2016; Topal et al., 2008b, a; Li and Zhou, 2018; Li et al., 2017), acoustic sensors (Golard and Talathi, 2021), magnetic sensors (Whitmire et al., 2016), Electrooculography (EOG) sensors (Bulling et al., 2008, 2009), or inertial measurement units (IMU) (Lanata et al., 2015) have been deployed on different kinds of wearables including glasses (Topal et al., 2008b, a; Zhang et al., 2011; Borsato and Morimoto, 2016; Ryan et al., 2008; Cho et al., 2012; Golard and Talathi, 2021; Li and Zhou, 2018; Mayberry et al., 2014, 2015), goggles (Bulling et al., 2008, 2009), hat (Lanata et al., 2015), and head-mounted devices (Whitmire et al., 2016; Noris et al., 2011; Kassner et al., 2014; Li et al., 2017). Among all these wearable eye tracking technologies, camera-based ones usually outperform others in terms of tracking performance and do not require lots of calibration data from new users. Many wearable eye trackers using cameras, especially on glasses, have become commercial and can be used as a reliable way to track eye movements continuously, such as Tobii Pro Glasses 3 (AB, 2022b), Pupil Labs (Invisible, Core, VR/AR add-ons) (Labs, 2022), Dikablis Glasses 3 (Ergoneers, 2022), and SMI Eye Tracking Glasses (iMotions, 2022). With these technologies, various novel gaze-based applications have been enabled, including detection of eye contacts (Ye et al., 2012), interaction with devices (Kytö et al., 2018; Matsubara et al., 2015; Paletta et al., 2014), monitoring mental health (Vidal et al., 2012; Li et al., 2020), etc.
@@
Despite of the promising tracking performance, current solutions to wearable eye tracking systems still have some limitations. First of all, many eye tracking systems above can only recognize discrete gestures (Zhang et al., 2011; Topal et al., 2008a, b; Bulling et al., 2008, 2009), limiting their performance in applications that need continuous tracking of the eyes. Camera-based wearable eye trackers can provide high accuracy in continuous eye tracking, but cameras are usually power-hungry, which makes them relatively impractical while deployed in wearables that need to be worn in everyday settings. To address this issue, (Mayberry et al., 2014) and (Mayberry et al., 2015) proposed low-power solutions to tracking gaze positions with cameras on glasses while maintaining promising accuracies. Despite of the impressive performance, changing lighting conditions can still be a problem for camera-based systems, as (Mayberry et al., 2015) demonstrated that the performance became worse in an outdoor setting. Besides, commercial eye trackers are usually expensive and do not provide open-source software for users, preventing them from being easily accessed and adapted by general users.
@@ 3.1.2. Acoustic Patterns for Continuous Eye Tracking
After receiving the reflected FMCW signals, we first apply a Butterworth band-pass filter with a cut-off frequency range of 18-21 kHz or 21.5-24.5 kHz on the signal to remove the signals in the frequency range that we are not interested in. It also helps protect the privacy of users because we remove the audible range of the signals. Then we further process the filtered signal to obtain unique acoustic patterns. According to prior research work (Wang et al., 2018; Li et al., 2022; Zhang et al., 2023b; Sun et al., 2023; Mahmud et al., 2023; Zhang et al., 2023a; Li et al., 2024), Echo Profile provides an accurate depiction of the status and movements of the reflecting objects in the environment. As a result, in this paper, we also use echo profiles as the acoustic patterns that our system monitors. As shown in Fig. 2 (a)-(c), echo profile is obtained by continuously calculating the cross-correlation between the received signals and transmitted signals. Fig. 1 demonstrates that different eye fixations and movements are correlated with different patterns in echo profiles. Based on these observations above, we believe that our GazeTrak system utilizing FMCW-based active acoustic sensing is able to track eye movements continuously with high accuracy.
@@ 3.2.2. Deep Learning Model
Because of the limited distance between the sensors on the glasses and the eyes, we are only interested in a certain range of the echo profiles (Fig. 2 (c)). As a result, we crop the echo profiles of each channel to get the center 70 pixels (23.8 cm) vertically. Then we randomly crop 60 pixels (20.4 cm) out of these 70 pixels for data augmentation purpose to make sure the system will not be severely impacted by the vertical shifting caused by remounting the device. To continuously track the gaze positions, we apply a sliding window of 0.3 seconds on the echo profiles. As a result, the dimension of the echo profile that we input into the deep learning model for one channel is 26 (0.3 s 50000 Hz 600 samples + 1) 60 (pixels). Because we use 2 speakers and 8 microphones in our system, which will be illustrated in Subsec. 4.2, we crop out the same dimension of echo profiles for all 2 8 = 16 channels, making the dimension of the input vector to the deep learning model as 26 60 16.
@@ 4.1. Hardware Design
In order to implement the FMCW-based active acoustic sensing technique mentioned in the section above, we chose Teensy 4.1 (PJRC, 2023b) as the micro-controller to provide reliable FMCW signal generation and receiving. We designed a PCB board to support two SGTL5000 chips which are the same as the one on the Teensy audio shield (PJRC, 2023a). With this customized PCB board plugged onto Teensy, it can support as many as 8 microphones and 2 speakers. We chose the speaker called OWR-05049T-38D (DigiKey, 2023) and the MEMS microphone called ICS-43434 (TDK, 2023) to support signal transmission and reception. We also built customized PCB boards for the speaker and the microphone to make them as small as possible. We used the Inter-IC Sound (I2S) buses on the Teensy 4.1 to transmit data between the Teensy 4.1 and the SGTL5000 chips, speakers and microphones. The collected data is stored in the SD card on Teensy 4.1. Fig. 3 (a)-(e) show these components.
@@ 5. User Study Procedure
We successfully recruited 20 participants (10 females and 10 males, 22 years old on average). Note that some participants participated in the study for multiple times to test different settings. The study was conducted in an experiment room on a university campus. During the study, the participants sit on a chair and put on the glasses form factor with our GazeTrak system. For each participant, we produced 12 sessions of instruction points. During the interval between sessions, participants were instructed to remove the device, place it on the table, and then put it back on. This step was taken to demonstrate that our system continued to function correctly even after the device was remounted. In each session, the instruction point moved to all the 100 pre-defined regions in a random order. The duration for which the instruction point remained at each position varied from 0.5 to 3.5 seconds, with an average of 2 seconds. As a result, the average length of each instruction session was 200 seconds. Before each session, there was a 15-second calibration process with the instruction point moving to the four corners of the screen and the center of the screen.
@@ 6.5.2. Real-world Noisy Environments
We trained a personalized model for each participant using 6 sessions of data collected in the quiet room. Then, the 2 testing sessions collected in each scenario were used to test the performance of our system in different environments. The average gaze tracking performance of our system across 10 participants remained consistent at and under two noisy environments, playing music and in the cafe, while the performance in the quiet room was . Overall, the average accuracy of gaze tracking did not change significantly with the presence of noise in the environment. We conducted a one-way repeated measures ANOVA test among the results of these three scenarios for all 10 participants and did not find a statistically significant difference (). This again validates that our system is not easily affected by environmental noise.
@@ 6.6. Impact of Different Glass Frames
We collected 12 sessions of data for each participant testing each glass frame. Since Subsec. 6.4 indicates that 6 sessions of training data is sufficient, we dumped the first 4 sessions for each glass frame and used the last 8 sessions to run a 4-fold cross validation in order to test the tracking performance of our system on different glasses. In this case, we can make sure that participants are familiar with the wearing of all the glass frames and eliminate the impact of some random factors. The evaluation result shows that the small glasses (F2) yielded a similar average performance to the original glasses (F1) (both at ), while the large glasses (F3) resulted in a relatively poorer average performance (at ), with a drop in performance of 15%. One possible reason for the performance difference is that the sensors on the larger glasses were much closer to the skin. Sometimes, the sensors may directly touch the skin, which could block the transmission and reception of signals, as we explained in Subsec. 4.2.
@@ 6.7.3. Usability
7. Data Preprocessing and Model Inference on MAX78002
In the previous evaluation, we recorded audio data with Teensy 4.1 first and run the signal processing and deep learning pipeline on a local server offline. To enable predictions of gaze positions in real-time on an MCU, we deployed the whole pipeline on a micro-controller with an ultra-low-power CNN accelerator (MAX78002 (Devices, 2022)).
@@ 7.2. Data Preprocessing
Before the deep learning model, we also need to apply a band-pass filter on the received signals and perform cross-correlation between received signals and transmitted signals to obtain echo profiles as described in Subsubsec. 3.1.2. In the implementation on MAX78002, we removed the band-pass filter since all the computations are done on the MCU and we do not have to use a band-pass filter to filter out the audible range of signals to protect users’ privacy. Besides, this helps reduce the processing time.
@@ 7.3. Accuracy and Refresh Rate
Given the limitation of the I2S interfaces on MAX78002, to test our system in a more realistic condition, we still use Teensy 4.1 to control the speakers and microphones and transfer the received audio data to MAX78002 via the serial port. This generates a steady stream of audio data to MAX78002 similar to the stream it would receive if it could accommodate 8 microphones by itself. In future, we will explore connecting microphones directly to MAX78002 using multi-channel audio protocols such as Time-division Multiplexing (TDM). Evaluation results showed that for ResNet-18 and MobileNet, MAX78002 spent 124.1 ms and 41.6 ms respectively loading the weights of the model. This is a one-time effort and can be done before running the real-time pipeline. Then it took 12 ms to load one instance and make an inference based on it for both ResNet-18 and MobileNet, giving a refresh rate of 83.3 Hz.
@@ 7.4. Power Consumption
If we can use MAX78002 to directly control speakers and microphones in future, we will be able to optimize the power efficiency and keep the overall power consumption of our real-time system around 95.4 mW, i.e., 79.0 mW (MAX78002 with ResNet-18 running at 30 HZ) + 16.4 mW (2 speakers and 8 microphones). We admit that this is just an estimate of the power of this real-time system and the power consumption of MAX78002 might increase if it does need to control the sensors but we do not expect it to be very high because the current power of MAX78002 already includes that of the CPU and the CNN accelerator running at full speed.
@@ 8.2.1. Head Movements
In the user study, we did not use a chin rest to fix the participants’ head so they could turn their head freely. However, we believe that a thorough evaluation on how head movements affect system performance is preferred in future work.
@@ Acknowledgements.
This work is supported by National Science Foundation (NSF) under Grant No. 2239569, NSF’s Innovation Corps (I-Corps) under Grant No. 2346817, the Ignite Program and the Ann S. Bowers College of Computing and Information Science at Cornell University. The authors would like to thank Prof. Susan Fussell and her lab for sharing the access to the commercial eye tracker Tobii Pro Fusion with us. We also appreciate the help of a Cornell student, Vipin Gunda, on the development of the pipeline on MAX78002.References
@@
- Li et al. (2024) Ke Li, Ruidong Zhang, Siyuan Chen, Boao Chen, Mose Sakashita, François Guimbretière, and Cheng Zhang. 2024. EyeEcho: Continuous and Low-power Facial Expression Tracking on Glasses. arXiv:2402.12388 [cs.HC]

### v2

@@ 1. Introduction
We conducted two rounds of user studies to evaluate the performance of GazeTrak. During the studies, each participant was asked to look at and follow the instruction points on the screen. In the first round of the study, 12 participants evaluated our first hardware prototype, where the microphones and speakers were glued on a glass frame. The average cross-session tracking accuracy was . It confirmed the optimal settings of the sensing system, which helped us design the final prototype. The final prototype (as shown in Fig. 1) features a more compact form factor, significantly lowers signal strength, and improves environmental sustainability as it can be attached to different glasses. To ensure consistent performance between the two prototypes, we conducted a second round of study with 10 participants, including some new participants, evaluating the final prototype. The final prototype achieved an average tracking accuracy of for cross-session scenarios and for in-session scenarios with a refreshing rate of 83.3 Hz. We made a demo video111https://youtu.be/XvNLNkfQY7Q to demonstrate the tracking performance and real-world applications of our system.
@@ 2. Related Work
In this section, we introduce the prior webcam-based, non-wearable, and wearable eye tracking systems.
@@ 2.3. Wearable Eye Tracking Technologies
To overcome the challenges that non-wearable eye tracking technologies face as described in the last two subsections, many wearable eye tracking technologies based on cameras (Zhang et al., 2011; Ryan et al., 2008; Noris et al., 2011; Cho et al., 2012; Kassner et al., 2014; Lanata et al., 2015; Mayberry et al., 2014, 2015), optical sensors (Borsato and Morimoto, 2016; Topal et al., 2008b, a; Li and Zhou, 2018; Li et al., 2017), acoustic sensors (Golard and Talathi, 2021), magnetic sensors (Whitmire et al., 2016), Electrooculography (EOG) sensors (Bulling et al., 2008, 2009), or inertial measurement units (IMU) (Lanata et al., 2015) have been deployed on different kinds of wearables including glasses (Topal et al., 2008b, a; Zhang et al., 2011; Borsato and Morimoto, 2016; Ryan et al., 2008; Cho et al., 2012; Golard and Talathi, 2021; Li and Zhou, 2018; Mayberry et al., 2014, 2015), goggles (Bulling et al., 2008, 2009), hat (Lanata et al., 2015), and head-mounted devices (Whitmire et al., 2016; Noris et al., 2011; Kassner et al., 2014; Li et al., 2017). Among all these wearable eye tracking technologies, camera-based ones usually outperform others in terms of tracking performance and do not require lots of calibration data from new users. Many wearable eye trackers using cameras, especially on glasses, have become commercial and can be used as a reliable way to track eye movements continuously, such as Tobii Pro Glasses 3 (AB, 2022b), Pupil Labs (Invisible, Core, VR/AR add-ons) (Labs, 2022), Dikablis Glasses 3 (Ergoneers, 2022), and SMI Eye Tracking Glasses (iMotions, 2022). With these technologies, various novel gaze-based applications have been enabled, including detection of eye contacts (Ye et al., 2012), interaction with devices (Kytö et al., 2018; Matsubara et al., 2015; Paletta et al., 2014), and monitoring mental health (Vidal et al., 2012; Li et al., 2020).
@@
Despite of the promising tracking performance, current solutions to wearable eye tracking systems still have some limitations. First of all, many eye tracking systems above can only recognize discrete gestures (Zhang et al., 2011; Topal et al., 2008a, b; Bulling et al., 2008, 2009), limiting their performance in applications that need continuous tracking of the eyes. Camera-based wearable eye trackers can provide high accuracy in continuous eye tracking, but cameras are usually power-hungry, which makes them relatively impractical while deployed in wearables that need to be worn in everyday settings. To address this issue, Mayberry et al. (Mayberry et al., 2014, 2015) proposed low-power solutions to tracking gaze positions with cameras on glasses while maintaining promising accuracies. Despite of the impressive performance, changing lighting conditions can still be a problem for these camera-based systems, as the performance became worse in an outdoor setting (Mayberry et al., 2015). Besides, commercial eye trackers are usually expensive and do not provide open-source software for users, preventing them from being easily accessed and adapted by general users.
@@ 3.1.2. Acoustic Patterns for Continuous Eye Tracking
After receiving the reflected FMCW signals, we first apply a Butterworth band-pass filter with a cut-off frequency range of 18-21 kHz or 21.5-24.5 kHz on the signal to remove the signals in the frequency range that we are not interested in. It also helps protect the privacy of users because we remove the audible range of the signals. Then we further process the filtered signal to obtain unique acoustic patterns. According to prior research work (Wang et al., 2018; Li et al., 2022; Zhang et al., 2023b; Sun et al., 2023; Mahmud et al., 2023; Zhang et al., 2023a; Li et al., 2024; Lee et al., 2024), Echo Profile provides an accurate depiction of the status and movements of the reflecting objects in the environment. As a result, in this paper, we also use echo profiles as the acoustic patterns that our system monitors. As shown in Fig. 2 (a)-(c), echo profile is obtained by continuously calculating the cross-correlation between the received signals and transmitted signals. Fig. 1 demonstrates that different eye fixations and movements are correlated with different patterns in echo profiles. Based on these observations above, we believe that our GazeTrak system utilizing FMCW-based active acoustic sensing is able to track eye movements continuously with high accuracy.
@@ 3.2.2. Deep Learning Model
Because of the limited distance between the sensors on the glasses and the eyes, we are only interested in a certain range of the echo profiles (Fig. 2 (c)). As a result, we crop the echo profiles of each channel to get the center 70 pixels (23.8 cm) vertically. Then we randomly select 60 consecutive pixels (20.4 cm) out of these 70 pixels for data augmentation purpose to make sure the system will not be severely impacted by the vertical shifting caused by remounting the device. To continuously track the gaze positions, we apply a sliding window of 0.3 seconds on the echo profiles. As a result, the dimension of the echo profile that we input into the deep learning model for one channel is 26 (0.3 s 50000 Hz 600 samples + 1) 60 (pixels). Because we use 2 speakers and 8 microphones in our system, which will be illustrated in Subsec. 4.2, we crop out the same dimension of echo profiles for all 2 8 = 16 channels, making the dimension of the input vector to the deep learning model as 26 60 16.
@@ 4.1. Hardware Design
In order to implement the FMCW-based active acoustic sensing technique mentioned in the section above, we chose Teensy 4.1 (PJRC, 2023b) as the micro-controller to provide reliable FMCW signal generation and receiving in multiple channels. We designed a PCB board to support two SGTL5000 chips which are the same as the one on the Teensy audio shield (PJRC, 2023a). With this customized PCB board plugged onto Teensy, it can support as many as 8 microphones and 2 speakers. We chose the speaker called OWR-05049T-38D (DigiKey, 2023) and the MEMS microphone called ICS-43434 (TDK, 2023) to support signal transmission and reception. We also built customized PCB boards for the speaker and the microphone to make them as small as possible. We used the Inter-IC Sound (I2S) buses on the Teensy 4.1 to transmit data between the Teensy 4.1 and the SGTL5000 chips, speakers and microphones. The collected data is stored in the SD card on Teensy 4.1. Fig. 3 (a)-(e) show these components.
@@ 5. User Study Procedure
We recruited 20 participants (10 females and 10 males, 22 years old on average). Note that some participants participated in the study for multiple times to test different settings. The study was conducted in an experiment room on a university campus. During the study, the participants sit on a chair and put on the glasses form factor with our GazeTrak system. For each participant, we produced 12 sessions of instruction points. During the interval between sessions, participants were instructed to remove the device, place it on the table, and then put it back on. This step was taken to demonstrate that our system continued to function correctly even after the device was remounted. In each session, the instruction point moved to all the 100 pre-defined regions in a random order. The duration for which the instruction point remained at each position varied from 0.5 to 3.5 seconds, with an average of 2 seconds. As a result, the average length of each instruction session was 200 seconds. Before each session, there was a 15-second calibration process with the instruction point moving to the four corners of the screen and the center of the screen.
@@ 6.5.2. Real-world Noisy Environments
We trained a personalized model for each participant using 6 sessions of data collected in the quiet room. Then, the 2 testing sessions collected in each scenario were used to test the performance of our system in different environments. The average gaze tracking performance of our system across 10 participants remained satisfactory at and under two noisy environments, playing music and in the cafe, while the performance in the quiet room was . Overall, the average accuracy of gaze tracking did not change significantly with the presence of noise in the environment. We conducted a one-way repeated measures ANOVA test among the results of these three scenarios for all 10 participants and did not find a statistically significant difference (). This again validates that our system is not easily affected by environmental noise.
@@ 6.6. Impact of Different Glass Frames
We collected 12 sessions of data for each participant testing each glass frame. Since Subsec. 6.4 indicates that 6 sessions of training data is sufficient, we discarded the first 4 sessions for each glass frame and used the last 8 sessions to run a 4-fold cross validation in order to test the tracking performance of our system on different glasses. In this case, we can make sure that participants are familiar with the wearing of all the glass frames and eliminate the impact of some random factors. The evaluation result shows that the small glasses (F2) yielded a similar average performance to the original glasses (F1) (both at ), while the large glasses (F3) resulted in a relatively poorer average performance (at ), with a drop in performance of 15%. One possible reason for the performance difference is that the sensors on the larger glasses were much closer to the skin. Sometimes, the sensors may directly touch the skin, which could block the transmission and reception of signals, as we explained in Subsec. 4.2.
@@ 6.7.3. Usability
7. Inference on MAX78002
In the previous evaluation, we recorded audio data with Teensy 4.1 first and run the signal processing and deep learning pipeline on a local server offline. To enable predictions of gaze positions in real-time on an MCU, we implemented the whole pipeline on a micro-controller with an ultra-low-power CNN accelerator (MAX78002 (Devices, 2022)).
@@ 7.2. Data Preprocessing
Before the deep learning model, we also need to apply a band-pass filter on the received signals and perform cross-correlation between received signals and transmitted signals to obtain echo profiles as described in Subsubsec. 3.1.2. In the implementation on MAX78002, to reduce processing time, we removed the band-pass filter since all the computations are done on the MCU and transmitting private data is no longer a concern.
@@ 7.3. Accuracy and Refresh Rate
Given the limitation of the I2S interfaces on MAX78002, to test our system in a more realistic condition, we still use Teensy 4.1 to control the speakers and microphones and transfer the received audio data to MAX78002 via the serial port. To accelerate the transmission speed, only the samples that are used for processing on MAX78002 are transferred. This generates a steady stream of audio data to MAX78002. In future, we will explore connecting microphones directly to MAX78002 using multi-channel audio protocols such as Time-division Multiplexing (TDM). Evaluation results showed that for ResNet-18 and MobileNet, MAX78002 spent 124.1 ms and 41.6 ms respectively loading the weights of the model. This is a one-time effort and can be done before running the real-time pipeline so it did not impact the refresh rate. Then it took 12 ms to load one instance and make an inference based on it in real-time for both ResNet-18 and MobileNet, giving a refresh rate of 83.3 Hz.
@@ 7.4. Power Consumption
If we can use MAX78002 to directly control speakers and microphones in future, we will be able to optimize the power efficiency and keep the overall power consumption of our real-time system around 95.4 mW, i.e., 79.0 mW (MAX78002 with ResNet-18 running at 30 HZ) + 16.4 mW (2 speakers and 8 microphones). One should keep in mind that this is just an estimate of the power of this real-time system and the power consumption of MAX78002 might increase if it does need to control the sensors but we do not expect it to be very high because the current power of MAX78002 already includes that of the CPU and the CNN accelerator running at full speed.
@@ 8.2.1. Head Movements
In the user study, we did not use a chin rest to fix the participants’ head so they could turn their head freely. However, we believe that how head movements affect the system performance should be evaluated in more details in the future.
@@ Acknowledgements.
This work is supported by National Science Foundation (NSF) under Grant No. 2239569, NSF’s Innovation Corps (I-Corps) under Grant No. 2346817, the IGNITE Innovation Acceleration Program, and the Ann S. Bowers College of Computing and Information Science at Cornell University. The authors would like to thank Prof. Susan Fussell and her lab for sharing the access to the commercial eye tracker Tobii Pro Fusion with us. We also appreciate the help of a Cornell student, Vipin Gunda, on the development of the pipeline on MAX78002.References
@@
- Lee et al. (2024) Chi-Jung Lee, Ruidong Zhang, Devansh Agarwal, Tianhong Catherine Yu, Vipin Gunda, Oliver Lopez, James Kim, Sicheng Yin, Boao Deng, Ke Li, Mose Sakashita, Francois Guimbretiere, and Cheng Zhang. 2024. EchoWrist: Continuous Hand Pose Tracking and Hand-Object Interaction Recognition Using Low-Power Active Acoustic Sensing On a Wristband. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, 21 pages. https://doi.org/10.1145/3613904.3642910
@@
- Li et al. (2024) Ke Li, Ruidong Zhang, Siyuan Chen, Boao Chen, Mose Sakashita, François Guimbretière, and Cheng Zhang. 2024. EyeEcho: Continuous and Low-power Facial Expression Tracking on Glasses. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, 24 pages. https://doi.org/10.1145/3613904.3642613